OpenStudy (anonymous):

The value of \[\sqrt{15-2\sqrt{15-2\sqrt{15-2\sqrt{15........}}}}\] is?

OpenStudy (jamesj):

If this has a value, call that value x. Then \[ x = \sqrt{15 - 2\sqrt{x}} \] Solve that equation.

OpenStudy (experimentx):

let \[ y = \sqrt{15-2\sqrt{15-2\sqrt{15-2\sqrt{15........}}}} \] square it \[ y^2 = 15-2\sqrt{15-2\sqrt{15-2\sqrt{15........}}} \] \[ y^2 = 15-2y \] solve for y

OpenStudy (jamesj):

oops yes, \[ x = \sqrt{15-2x} \]

OpenStudy (anonymous):

y=3 or y=-5

OpenStudy (anonymous):

@experimentX then

OpenStudy (anonymous):

is it 3 .....

OpenStudy (experimentx):

sorry ... from JanesJ's method ... it's 3

OpenStudy (experimentx):

you are not getting negative no from \( \sqrt{ ....}\)

OpenStudy (anonymous):